OpenStudy (anonymous):

I have to draw my question...click on this to see the math question, please!!

OpenStudy (anonymous):



OpenStudy (anonymous):

one angle is 68 degrees

OpenStudy (anonymous):

the hypotenus is 9

OpenStudy (anonymous):

you have to solve the right triangle, and round each side the nearest hundreth.

OpenStudy (lgbasallote):

sounds like a job for sines and cosines huh

OpenStudy (anonymous):

would i use a^2 + b^2=c^2?

OpenStudy (anonymous):

First, J is 180-90-68=22 degrees. Now we're going to use the law of sines, with "J,K,L" as angles and "j,k,l" as their opposite sides: \[\begin{align} &\frac{\sin K}{k}=\frac{\sin J}{j}\\ &=\frac{1}{9}=\frac{.37}{j}\\ j= .04 \end{align}\] See if you can do L the same way.

OpenStudy (lgbasallote):

you use \[\sin \theta = \frac{opposite}{hypotenuse}\] and \[\cos \theta = \frac{adjacent}{hypotenuse}\]

OpenStudy (anonymous):

Except that I typed that into my calculator wrong. j is 3.37.

OpenStudy (anonymous):

i know the hypotenus is 9. but the opposite is a variable, k. i don't really know what sin and cos means

OpenStudy (anonymous):

are u using a graphing calculator?

OpenStudy (lgbasallote):

it's in your calculator... sin and cos

OpenStudy (anonymous):

k

OpenStudy (anonymous):

hello? um, can you give me a moment while i figure this out?

OpenStudy (anonymous):

i think i need a bit more help, please

OpenStudy (anonymous):

OK, where are you stuck?

OpenStudy (anonymous):

i understand how sink/k = sinJ/j, but how did it equal 1/9?

OpenStudy (anonymous):

and then, how did you get .37/j? i'm new to the graphing calculator stuff...

OpenStudy (anonymous):

This isn't a graphic calculator thing, this is really a trigonometry issue. Have you learned sin, cos, and tan yet?

OpenStudy (anonymous):

not yet, i'm just starting geometry this fall

OpenStudy (anonymous):

There's no way I know of to get the other sides without trigonometric functions...anyone else?

OpenStudy (anonymous):

could you show me the trigonometry way? i have a graphing calculator

OpenStudy (anonymous):

You should really understand what those trig functions mean first. So here we go:

OpenStudy (anonymous):

k

OpenStudy (anonymous):